OpenStudy (anonymous):

Let f(x) = \begin{cases} -2 x+b, &\text{if}\ x< 5\\ \frac{-150}{x-b}, &\text{if}\ x\geq5 \end{cases} There are exactly two values for b which make f(x) a continuous function at x=5. The one with the greater absolute value is

OpenStudy (freckles):

set left limit equal to right limit then solve for b

OpenStudy (anonymous):

replace \(x\) by \(5\) in both expressions, set them equal, solve for \(b\) you will get a quadratic equation in \(b\) hence the two solutions

OpenStudy (anonymous):

which is another somewhat more prosaic was of exactly what @freckles said

OpenStudy (anonymous):

i get -10+b=150/5-b and dont know what to do after

OpenStudy (anonymous):

start with \[(-10+b)(5-b)=150\] and solve that one

OpenStudy (anonymous):

yes but then you get -bsquared+15b-300

myininaya (myininaya):

is -50-150=-300?

OpenStudy (anonymous):

ahh pellett thanks!!!

OpenStudy (anonymous):

it still doesnt make sense

OpenStudy (anonymous):

unless i distribute wrong

myininaya (myininaya):

it really looks like your answer will be ugly unfortunately

myininaya (myininaya):

oh you know what

myininaya (myininaya):

i think there was a -150/(x-b)

myininaya (myininaya):

was that considered earlier

myininaya (myininaya):

not 150/(x-b)

OpenStudy (anonymous):

ohh i put put postive instead of negative 150

OpenStudy (anonymous):

i got it thank you very much

myininaya (myininaya):

that should have a better pretty answer :)

OpenStudy (anonymous):

its 20

myininaya (myininaya):

gj you!